OpenStudy (anonymous):

Find the area of one segment formed by a square with sides of 6" inscribed in a circle

OpenStudy (anonymous):

This is the diagram, by Pythagorus' theorem, the diameter of the circle is 6√2

OpenStudy (anonymous):

r = 3√2 Area of the entire circle = pi * r^2 = pi * 18 Area of square = 6*6 = 36 Area of 4 sectors = Area of circle - square = 18pi - 36 Area of 1 sector = (18pi - 36)/4 = 4.5pi - 9

OpenStudy (anonymous):

Of course, stick your units and approximate if you must